OpenStudy (anonymous):

the ratio of the diameter of the conical tent to its height is 3 to 5. if the height is 10ft. find the amount of air inside.

Directrix (directrix):

the ratio of the diameter of the conical tent to its height is 3 to 5. the height is 10ft. -------------------------- This means d/h = 3/5 d/10 = 3/5 You solve for d and post the result here. @LOL10